# ControlMM: Controllable Masked Motion Generation

### v1

@@ A.7 Analysis of Logits Editing and Motion Control Model
Without Motion Control Model
In the first iteration (iteration 0), the probability is low but increases significantly in the subsequent iterations. After applying Logits Editing, the probability improves slightly, as shown in Fig. 8 and 7. Eventually, the probability saturates in the later iterations (see Figure 11). Since the probability of most token predictions approaches one, Logits Editing cannot further modify the logits, preventing any updates to the trajectory.
Figure 7: The maximum probability of the each token without Motion Control Model before Logits Editing of each all 49 tokens and 10 iterations.
Figure 8: The maximum probability of the each token without Motion Control Model after Logits Editing of each all 49 tokens and 10 iterations.
With Motion Control Model
With the introduction of the Motion Control Model, the probability of token predictions is significantly higher in the initial iteration compared to the scenario without the Motion Control Model, as illustrated in Figures 9 and 10. Moreover, the maximum probability does not saturate to one, indicating that there is still room to adjust the logits for trajectory editing.
This enhancement leads to improved generation quality within fewer iterations, as detailed in Section A.6. Notably, just 4 iterations using the Motion Control Model yield a quality comparable to that achieved in 10 iterations without it, where the latter still exhibits suboptimal quality and high average error.
Figure 9: The maximum probability of the each token with Motion Control Model before Logits Editing of each all 49 tokens and 10 iterations.
Figure 10: The maximum probability of the each token with Motion Control Model after Logits Editing of each all 49 tokens and 10 iterations.
Average of maximum probability of all tokens in each iteration To clearly illustrate the increasing probability or confidence of the model predictions across all 10 iterations, as shown in Fig. 11. In this figure, the blue line represents the average probability of token predictions With the Motion Control Model, while the red line denotes the average probability Without the Motion Control Model. The solid line indicates the average probability prior to the application of Logits Editing. This shows that the probability increases significantly in the very first iteration for the With the Motion Control Model.
@@
A.8 Approximated Mask Embedding for Decoder
Motion Control Model requires the relative positions between the predicted motion from the ongoing unmasking process, which requires [Mask] tokens for an initial motion token generation. However, [Mask] tokens are only used for the Masked Transformer, and there is no [Mask] token in the codebook, making it impossible to reconstruct motion from Masked Transformer embeddings. To address this issue, we approximate the [Mask] token for the codebook space.
In this section, we visualize the embedding of the [Mask] token (black) compared to all Transformer tokens (red), as shown in Fig. 12. The visualization indicates that the [Mask] embedding is approximately the average of all embeddings. By using the average of all embeddings for the [Mask] position, we can utilize the relative differences between the generated motion and the control motion for Motion Control Model.
Figure 12: t-SNE visualization of the embeddings for all Transformer tokens (red), comparing to the [Mask] token (black).
A.9 Body Part Timeline Control
@@
It is important to note that this approach may lead to out-of-distribution generation since the model has not been trained on combinations of multiple body parts with their associated text prompts. However, our model handles out-of-distribution generation effectively due to the use of Logits Editing and Codebook Editing.
@@
A.10 Dual-Space Categorical Straight-Through Estimator
To reconstruct motion from Transformer tokens, the tokens must first be mapped to their corresponding codebook embeddings using the same indices before being fed into the decoder. However, this index-based lookup operation is inherently non-differentiable, which obstructs guidance from the generated motion through the gradient backpropagation.
Dual-Space Categorical Straight-Through Estimator (DCSE) performs weighted average sampling of the codebook $C$ w.r.t. the probability distribution $p$. Given the output logits $l$ from the Transformer, instead of using the non-differentiable $\arg\max$ operation to select embedding from the codebook, we apply the Gumbel-Softmax function (Jang et al., 2017) to obtain a probability distribution as a smooth differentiable approximation alternative to the $\arg\max$ operation, producing $k$-dimensional sample vectors $p$.
$p_{i}=\frac{\exp\left((\ell_{i}+g_{i})/\tau\right)}{\sum_{j=1}^{k}\exp\left(\ell_{j}/\tau\right)}$ (10)
where $\tau$ refers to temperature and $g$ represents Gumbel noise with $g_{1},\dots,g_{k}$ being independent and identically distributed (i.i.d.) samples from a $\text{Gumbel}(0,1)$ distribution. The $\text{Gumbel}(0,1)$ distribution can be sampled via inverse transform sampling by first drawing $u\sim\text{Uniform}(0,1)$ and then computing $g=-\log(-\log(u))$.
From sample vectors $p$, the approximated embedding can be obtained from weighted sampling of Transformer token space $e_{j}$
$e_{t}=\sum_{j=1}^{k}p_{i}\cdot e_{j}$ (11)
or from code $c_{j}$ in Codebook $C$.
$e_{c}=\sum_{j=1}^{k}p_{i}\cdot c_{j}$ (12)
In our implementation, we adopt the configuration from MoMask (Guo et al., 2023), using a codebook embedding size of 512 and a Transformer token size of 384. With this setup, we demonstrate that conversion across different spaces is feasible, even when the embedding sizes differ, as long as both spaces refer to the same set of indices. This allows for flexible representation across latent spaces while maintaining consistency in how the embeddings are referenced.
Figure 14: (a) Dual-Categorical Straight-Through Estimator (b) Mask Motion Control Model
A.11 Cross Combination
We follow the evaluation Cross Combination from OmniControl (Xie et al., 2023), evaluating multiple combinations of joints as outlined in Table 1. A total of 63 combinations are randomly sampled during the evaluation process as follow.

### v2

@@ A.10 Analysis of Logits Optimization and Logits Regularizer
Without Logits Regularizer
In the first step (step 0), the probability is low but increases significantly in the subsequent steps. After applying Logits Optimization, the probability improves slightly, as shown in Fig. 6 and 7. Eventually, the probability saturates in the later steps (see Figure 10). Since the probability of most token predictions approaches one, Logits Optimization cannot further modify the logits, preventing any updates to the trajectory.
Figure 6: The maximum probability of the each token without Logits Regularizer before Logits Optimization of each all 49 tokens and 10 steps.
Figure 7: The maximum probability of the each token without Logits Regularizer after Logits Optimization of each all 49 tokens and 10 steps.
With Logits Regularizer
With the introduction of the Logits Regularizer, the probability of token predictions is significantly higher in the initial step compared to the scenario without the Logits Regularizer, as illustrated in Figures 8 and 9. Moreover, the maximum probability does not saturate to one, indicating that there is still room to adjust the logits for trajectory editing.
This enhancement leads to improved generation quality within fewer steps, as detailed in Section A.9. Notably, just 4 steps using the Logits Regularizer yield a quality comparable to that achieved in 10 steps without it, where the latter still exhibits suboptimal quality and high average error.
Figure 8: The maximum probability of the each token with Logits Regularizer before Logits Optimization of each all 49 tokens and 10 steps.
Figure 9: The maximum probability of the each token with Logits Regularizer after Logits Optimization of each all 49 tokens and 10 steps.
Average of maximum probability of all tokens in each step To clearly illustrate the increasing probability or confidence of the model predictions across all 10 steps, as shown in Fig. 10. In this figure, the blue line represents the average probability of token predictions With the Logits Regularizer, while the red line denotes the average probability Without the Logits Regularizer. The solid line indicates the average probability prior to the application of Logits Optimization. This shows that the probability increases significantly in the very first step for the With the Logits Regularizer.
@@
A.11 Ambiguity of Motion Control Signal
In the image domain, pixel control signals can be directly applied, and uncontrolled regions are simply zeroed out. However, for motion control, zero-valued 3D joint coordinates are ambiguous: they may indicate that a joint is controlled with its target position at the origin in Euclidean space, or that the joint is uncontrolled. To resolve this ambiguity, we concatenate the joint control signal with the relative difference between the control signal and the generated motion, forming the final joint control guidance $s$.
Figure 11: The difference between control signals: (a) Image Control: 0 means no control, 1 means control. (b) Absolute Motion Control: ambiguous between control signal at origin and no control. (c) Relative Motion Control: no ambiguity. Black curve: spatial control signal. Blue curve: decoded spatial signal from generated motion
A.12 Body Part Timeline Control
@@
It is important to note that this approach may lead to out-of-distribution generation since the model has not been trained on combinations of multiple body parts with their associated text prompts. However, our model handles out-of-distribution generation effectively due to the use of Logits Optimization.
@@
A.13 Cross Combination
We follow the evaluation Cross Combination from OmniControl [55], evaluating multiple combinations of joints as outlined in Table 2. A total of 63 combinations are randomly sampled during the evaluation process as follow.